You will find the size and field of action of the resultant of two parallel forces with sizes 40N and 60N, whose impactors are at a distance of 2m. The forces act: a, In the same direction b. In the opposite direction


1
Expert's answer
2021-06-14T08:18:27-0400

The resultant of the system if the forces are in the same direction

"R=\sum F = 40 N+ 60N=100N"


Assume distance d from one end of the system say point A

"\sum M_A^F = \sum F \times x"

"\sum M_A^F = 100 \times 2 = 200N.m"

"M_A^R=\sum M_A^F"

"100 d=200 \implies d= 2 m"

Thus R =100 N , 2 m to the right


The resultant of the system if the forces are in a different direction

"R=\sum F = -40 N+ 60N=20N"


Assume distance d from one end of the system say point A

"\sum M_A^F = \sum F \times x"

"\sum M_A^F = 20 \times 2 = 40N.m"

"M_A^R=\sum M_A^F"

"20 d=40 \implies d= 2 m"

Thus R =20 N , 2 m to the right